OpenStudy (anonymous):

log subscript 3x+2 of 125=3 HELPPPPPPPPPPPP ASAP

OpenStudy (kirbykirby):

\[ \large \log_{3x+2}125=3\\ \large \implies (3x+2)^3=125\]

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

PLEASE SIR HELP ME

OpenStudy (anonymous):

I NEED THIS ANSWER SO BADD

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@study100

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@wyattp17

OpenStudy (anonymous):

@Fire_Fighter_Of_Miami

OpenStudy (kirbykirby):

\[\large (3x+2)^3=(3x+2)^2(3x+2)=125\\ \large =(9x^2+12 x + 4)(3x+2)=125\\ \large =27x^3+18x^2+36x^2+24x+12x+8=125\\ \large =27x^3+54x^2+36x-117=0 \\ \large =9(3x^3+6x^2+4x-13)=0\\ \\\text{using the factor theorem, 1 is a zero of the polynomial in ( ), so (x-1) is a factor} \\ \text{then use long division to divide} 3x^3+6x^2+4x-13 \text{ into }x-1 \\ = \large 9\left[(x-1)(3x^2+9x-13) \right]=0 \\ \] This means that \(x-1=0\) or \(3x^2+9x-13=0\) So you can solve the 2nd equation using the quadratic formula

OpenStudy (anonymous):

I want to thank you but its too late i hope your efforts arent in vain though

OpenStudy (anonymous):

can someone help me solve the 2nd equation

OpenStudy (anonymous):

I really need help im in a bad jam right now

OpenStudy (kirbykirby):

\[x = \frac{-9\pm\sqrt{9^2-4(3)(-13)}}{2(3)}=\frac{-9\pm 5i\sqrt{3}}{6} \]

OpenStudy (kirbykirby):

hm actually I just noticed my typo.. it should be \(+13\) not \(-13\)

OpenStudy (anonymous):

I need to be 100% positive this is right i cant afford to screw up again i take my dba again at 1 will you still be online ?

OpenStudy (kirbykirby):

I should be

OpenStudy (anonymous):

could you help me when i take it it would be so greatly appreciated im just trying to get out of this class the math they give us is ridiculous and i have constant ridicule from my teacher and i honestly want to learn but not in this hostile enviroment

OpenStudy (kirbykirby):

ok i can try

OpenStudy (kirbykirby):

what class is this for?

OpenStudy (anonymous):

Alg 2

OpenStudy (anonymous):

i will call her back at 105 i need your help im going to review some more also

OpenStudy (anonymous):

she expected me to solve that problem in like 2 minutes

OpenStudy (kirbykirby):

Maybe trial and error? By inspection, x = 1 is a solution (the other 2 solutions are complex.. maybe you didn't cover those yet?)

OpenStudy (anonymous):

the answers i gave her were x=125 x=58 also x=37 all wrong

OpenStudy (anonymous):

I just need to get past this because the rest of the class are subjects i know but im stuck

OpenStudy (kirbykirby):

ok

OpenStudy (anonymous):

do you know anyone else that can help

OpenStudy (anonymous):

Ok i will call her back you ready ?

OpenStudy (kirbykirby):

ok

OpenStudy (anonymous):

Im going to call her now

OpenStudy (anonymous):

this is bullpellet she says i have to score higher on my test nvm

OpenStudy (kirbykirby):

o.o ok

OpenStudy (anonymous):

shes making me score high on a practice test or i cant continue in the course im sure thats breaking some type of rule

OpenStudy (kirbykirby):

:S I have no idea though... I'm not even in the U.S. so I don't even know how it works there